OpenStudy (anonymous):

just a simple fraction question! 5 1/6 ÷ 5 2/3 i already did it, but i'm not sure if it's right. i got 2 3/14? idk it's the only one of my answers i'm not sure of..

OpenStudy (anonymous):

31/34

OpenStudy (anonymous):

ya.. that's what this calculator thing i used kept saying. but i don't understand how to get that answer? i'll try again

OpenStudy (anonymous):

It also says 0.91176470588235294, or 0.9117647059

OpenStudy (anonymous):

\[5\frac{ 1 }{ 6 } \div 5\frac{ 2 }{ 3 }\]

OpenStudy (anonymous):

lol ik i'm pretty sure a decimal is not a valid answer though

OpenStudy (anonymous):

Answer: = 31/6 ÷ 17/3 = (31 × 3) / (17 × 6) = 93/102 = 31/34

OpenStudy (anonymous):

oh my god. i finally realized what i did wrong. i kept writing 7 instead of 17. oops! anyway thanks so much for your help :)